OpenStudy (anonymous):

EVALUATE AND IF POSSIBLE SIMPLIFY -X-13/26 x=-13

OpenStudy (anonymous):

-X-13/26 x=-13 if its like this (-X-13)/26 subs x=-13 = (-13-13)/26 =-26/26 =1 ans,.. but if its like this -X-(13/26) x=-13 subs x=-13 =-13-(13/26) =-13-(1/2) =-27/2 ans.